OpenStudy (anonymous):

solve for C a=1/3(a+b+c)

OpenStudy (mertsj):

1. Multiply both sides by 3 2. Subtract a from both sides 3. Subtract b from both sides.

OpenStudy (anonymous):

A. c = 3A – (a – b) B. c = 3A – a – b C. c = 3a + 3b + A D. c = 3Aa

OpenStudy (anonymous):

so 3a=(a+b+c)

OpenStudy (mertsj):

Do you have A and a or are they both a ?

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

I was wondering that.

OpenStudy (anonymous):

3=(b+c)

OpenStudy (anonymous):

then.. 3c?

OpenStudy (anonymous):

idk about the a's

OpenStudy (mertsj):

Look at the problem you were given. Does it have A in it?

OpenStudy (anonymous):

yes, A=1/3(a+b+c)

OpenStudy (mertsj):

Then: 1. Multiply both sides by 3 2. Subtract a from both sides 3. Subtract b from both sides.

OpenStudy (anonymous):

1. 3A=1/3(a+b+c) 2. 3= 1(b+c) 3. b-3=c

OpenStudy (mertsj):

1. You multiplied the left side by 3 but not the right side.

OpenStudy (anonymous):

1. 3A=1/9(a+b+c)

OpenStudy (mertsj):

What is: \[3\times \frac{1}{3}=\]

OpenStudy (anonymous):

1. 3A= 3/9(a+b+c)

OpenStudy (mertsj):

\[3 \times \frac{1}{3}= \]

OpenStudy (anonymous):

3/9

OpenStudy (mertsj):

Would you please do that problem?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

cross divide

OpenStudy (anonymous):

3/1 x 1/3 = 1/1 x1/1

OpenStudy (mertsj):

So what is: \[3 \times \frac{1}{3}(a+b+c)=\]

OpenStudy (anonymous):

1(a+b+c)

OpenStudy (mertsj):

And what is 1(a+b+c)

OpenStudy (anonymous):

a+b+c

OpenStudy (mertsj):

Now proceed with the three steps of the problem I outlined above.

OpenStudy (anonymous):

so then 3A=a+b+c 2. 3A=b+c 3. 3A=c

OpenStudy (mertsj):

1. Multiply both sides by 3 2. Subtract a from both sides 3. Subtract b from both sides.

OpenStudy (mertsj):

2. Subtract a FROM BOTH SIDES!!!!!!

OpenStudy (anonymous):

2. 3= b+c 3. b-3=c

OpenStudy (anonymous):

my answer isn't even close to one of the choices

OpenStudy (anonymous):

A. c = 3A – (a – b) B. c = 3A – a – b C. c = 3a + 3b + A D. c = 3Aa